5. Change the logarithmic expression to an equivalent expression involving an exponent.
log[2](1/32)=k
can someone help me out?

Answer by lwsshak3(11628)

5. Change the logarithmic expression to an equivalent expression involving an exponent.
log[2](1/32)=k
**
log[2](1/32)=k
log2(1/2^5)=k
log2(2^-5)=k
convert to exponential form:(base(2) raised to log of number(k)=number(2^-5)
2^k=2^-5
k=-5